OpenStudy (anonymous):

Can someone help me do this with complete the square? -x^2+y^2-18 x-14 y-132 = 0

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

i believe i have some of it done actually -(x2-18x)+81 +(y2-14y)49=132+81+49 ?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf -x^2+y^2-18x-14y-132 = 0\\ \quad \\\implies -(x^2+18x)+(y^2-14y)-132 = 0 \\ \quad \\ -(x^2+18x+{\color{red}{ 9}}^2)+(y^2-14y+{\color{red}{ 7}}^2)-132-{\color{red}{ 9}}^2-{\color{red}{ 7}}^2 = 0\)

OpenStudy (jdoe0001):

\(\bf -x^2+y^2-18x-14y-132 = 0 \\ \quad \\ \implies -(x^2+18x)+(y^2-14y)-132 = 0 \\ \quad \\ -(x^2+18x+{\color{red}{ 9}}^2)+(y^2-14y+{\color{red}{ 7}}^2)-132-{\color{red}{ 9}}^2-{\color{red}{ 7}}^2 = 0 \\ \quad \\ -(x+9)^2+(y-7)^2=262\implies (y-7)^2-(x+9)^2=262\)

OpenStudy (anonymous):

so after that just 262/262 and distribute it to the other side ?

OpenStudy (jdoe0001):

if you want to get the hyperbola equation, yes

OpenStudy (anonymous):

I always get confused on where to do b/2^2 xD

OpenStudy (anonymous):

wwould it be better to square root the 262 on the left side? or leave it as is

OpenStudy (jdoe0001):

hmm possible just leave it as it's

OpenStudy (anonymous):

Alright thanks!

OpenStudy (jdoe0001):

yw